# Is Limited Information Enough? An Approximate Multi-agent Coverage Control in Non-Convex Discrete Environments

### 08 Jan 2024 v1

@@ 3.1. High Level Description of the Algoritm
Note that $\mathcal{N}^{2}_{i}$ is introduced for the scalability of the algorithm and is not necessary for the results above. Theorem 5 and 7 hold with the naive definition of neighborhood $\mathcal{N}_{i}$. Note that when there are multiple candidates of $i_{\min}$, Algorithm 1 forms a forest that consists of disjoint trees rooted at those candidates, and the problem is split into independent subproblems for each communication tree. To ensure the theorems hold, we replace function Neighbor in Line 6 in Algorithm 2 with Algorithm 3 to synchronize the timing of updating the communication trees. The algorithm introduces the flag $s_{i}$ for synchronization. Agent $i$ is initialized with $s_{i}=False$ and updates it as $s_{i}=True$ when $i$ is picked up at Line 5 in Algorithm 2. With this flag, Algorithm 3 checks whether all the agents have been processed (Line 2). If yes, the algorithm resets the memory (Line 3) and tries to update the communication trees of all agents (Line 5). Otherwise, the algorithm tries to update each communication tree independently. To this end, the set of agents in the tree is identified (Line 7) and the flags of those agents are initialized (Line 8-9). $\textsc{TreeMember}(i)$ is a function to identify the member of the communication tree to which $i$ belongs. Then the procedure CommTree of $i$ is triggered to construct the communication trees (Line 10) and the algorithm returns the new neighborhood $\mathcal{N}^{2}_{i}$. This synchronization ensures that the entire communication forest is updated only when all the trees are processed once, thereby ensuring the progress of algorithm convergence.
Algorithm 3 Synchronized update of neighborhood
1:procedure $\mathcal{N}^{2}_{i}=$Neighbor($\mathcal{P},i$)
2:     if $s_{k}=True,\forall k\in\mathcal{N}$  then
3:         $s_{k}=False,\forall k\in\mathcal{N}$
4:     if $s_{k}=False,\forall k\in\mathcal{N}$ then
5:         $\hat{u}_{k}=u_{k},D_{k}=-1,\mathcal{\hat{N}}_{k}=\mathcal{N}_{k},\forall k\in\mathcal{N}$
6:     else
7:         $\mathcal{N}^{tree}_{i}=\textsc{TreeMember}(i)$
8:         $\hat{u}_{k}=u_{k},D_{k}=-1,\forall k\in\mathcal{N}^{tree}_{i}$
9:         $\mathcal{\hat{N}}_{k}=\mathcal{N}_{k}\cap\mathcal{N}^{tree}_{i},\forall k\in\mathcal{N}^{tree}_{i}$
10:     $D_{i}=\textsc{CommTree}(i)$
11:     $\mathcal{N}^{2}_{i}=\{i,D_{i}\}$
12:     Return $\mathcal{N}^{2}_{i}$
3.2. Complexity Analysis
As mentioned previously, Algorithm 2 requires solving as a subroutine an NP-hard problem (to compute the neighborhood optimum $M_{n_{i}}$ in Line 10 of Algorithm 2). This computation is the bottleneck of the algorithm, and its computational complexity is $\mathcal{O}(|\mathcal{P}_{n_{i}}|^{n_{i}})$: we need to compute for an agent the optimal way to allocate her neighbors within her partition $\mathcal{P}_{n_{i}}$. In the general case, the size of the neighborhood $n_{i}$ can be as high as $n-1$. This occurs when an agent is at the center of a circle and can be neighbor of all the agents located in that circle. This makes the algorithm non-scalable. Nevertheless, we introduce the neighborhood with the constant size ($n_{i}=2$), which makes the algorithm polynomial in $n$.
@@ 3.2. Complexity Analysis
As for the communication complexity of the algorithm, we start by bounding the convergence rate, which is the number of iterations before convergence. Let $d_{\max}$ be the upper bound of the distance between any two points in the environment $C$, and $\Delta v$ be the resolution limit of the weight $v_{c}$. In the case of discrete settings, the potential function $\phi(x,\mathcal{P})$ consists of the elemental term $v_{c}\leavevmode\nobreak\ g(|x_{i}-c|)$ and then the improvement in the potential function for each iteration is lower bounded by $\epsilon=\Delta v\cdot g(d_{\max})$. (In the case of continuous settings, we can regard $\epsilon$ as the agents’ resolution limit of utility). Note that the convergence of the algorithm is guaranteed by Theorem 5. For each iteration, Algorithm 2 picks an agent one by one and checks if the potential function can be improved or not. Before convergence, at least one out of $n$ agents improves the potential function. Then the convergence rate $\alpha$ is bounded as $\alpha\leq n[\phi(x^{*},\mathcal{P}^{*})-\phi(x,\mathcal{P})]/\epsilon$, where $(x,\mathcal{P})$ and $(x^{*},\mathcal{P}^{*})$ are the initial allocation and the allocation after convergence, respectively.
For each iteration, the agents synchronize the start and the end timing of the step, share the global information, share their private information to compute a neighborhood optimum, and finally share the neighborhood optimum solution. Agents also need to share the timing of updating the communication trees in Algorithm 3. To this end, we define additional global variables $s_{True}$ and $s_{False}$ that $s_{True}=True$ if $s_{i}=True$ for all agents, $s_{False}=True$ if $s_{i}=False$ for all agents, otherwise they are $False$. The communications to share the global information require at most $\mathcal{O}(n^{2})$ messages. The message to share the global information $(u_{\min},i_{\max}^{+},x_{i_{\min}},s_{True},s_{False})$ has a fixed length regardless of the problem size. If we assume $1/g$ is a polynomial function, $1/\epsilon$ and $\alpha$ are also polynomial in the problem size, and then the communication complexity is also polynomial.
@@ 4. Experiments
In order to validate the practical efficiency and scalability of our approach, we ran simulations. First, we evaluate the efficiency with small graphs, then we evaluate the scalability with larger graphs.
@@
The neighborhood optimal approach proposed in Section 3.1, coined in the following as NBO, is compared to:
@@
  • (SOTA) the algorithm of Sadeghi et al. Sadeghi et al. (2020). In a nutshell, for a given agent $i$, the algorithm first tries to perform individual moves to maximize the social welfare of her neighborhood $\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x)$. If no such move is improving, it considers coordinated moves with a single other agent $j$, in the sense that $i$ would move and $j$ would take the place of agent $i$. The algorithm first considers neighbors, and then (via neighborhood communication), may consider agents further away. However, these coordinated moves only involve two agents at most.
  • (CGR) the centralized greedy algorithm that allocates agents one by one starting from the environment with nobody.
We evaluate the performance of the algorithms above with different shapes of the environments (Figure 6). In addition of these shapes, we also use the small bridge setting and OR library dataset shown in Yun and Rus (2014). More details of the shapes are in the supplementary material.
@@ 4.1. Efficiency
First, we evaluate the efficiency of the proposed neighborhood optimum algorithm, by comparing it with an optimal solution and the benchmarks. The efficiency is measured with the ratio $G(x)/G(x^{*})$ where $x$ is a solution of the algorithm and $x^{*}$ is an optimal solution. Since finding an optimal solution is NP-hard, the simulation is conducted on a 1D chain where $|C|=20,|\mathcal{C}_{+}|=10$, and $N=5$.
In that case, the proposed algorithm outperforms both benchmarks and improves the efficiency by about 21.7% points compared to SOTA (Figure 7). The efficiency ratio of NBO is 90% which is much better than the theoretical approximation ratio of 2. We examine the detail with an illustrative example in the supplementary material.
Due to the NP-hardness, it is difficult to compare with optimal solutions in larger settings. Instead, we compute the efficiency ratio with the centralized greedy (CGR) as a basis. Table 1 summarizes all the results obtained. The quantitative comparison of the efficiency shows similar results as in Figure 7 that the proposed method outperforms the benchmarks in all cases. Note that NBO even outperforms CGR sometimes despite the following two disadvantages of NBO compared to CGR: 1. NBO uses only limited information while CGR uses full information, 2. NBO’s initial state is less favorable compared to CGR’s matroidal setting.
@@
Table 1. Comparison to VVP performance: Efficiency ratio $G(x)/G(x^{CGR})$ where $x$ is a solution of each algorithm and $x^{CGR}$ is a solution of CGR. Each cell shows mean ± standard deviation. Larger values mean better results.
@@ 4.2. Scalability
We evaluate the scalability of the proposed algorithm by changing the size of the space $|C|$ and the number of agents $n$. The results show that:
  • The runtime grows with a polynomial order of $|C|$.
  • The runtime decreases when $n$ increases because the size of partitions $|\mathcal{P}_{n_{i}}|$ also decreases.
This result closely aligns with the theoretical complexity $\mathcal{O}(|\mathcal{P}_{n_{i}}|^{2})$ and demonstrates the applicability of the proposed method to real-world large-scale problems. Due to the space limit, we defer the details to the supplementary material.
@@ 5. Conclusion
To guarantee the performance of distributed coverage control, we extend the approach of Marden (2016) to general dimensional settings for discrete environments, by communicating additional information on the position of the minimum utility agent. The algorithm guarantees convergence to a neighborhood optimum solution, even in challenging non-convex settings and achieves the approximation ratio of 2, which is equivalent to the case of full information on all agents. This approximation ratio surpasses that of the state-of-the-art method presented in Sadeghi et al. (2020). While both the approach of Sadeghi et al. (2020) and ours are more communication demanding than simple best responses based on Voronoi partitioning, this remains manageable and prioritizes local interactions.
Furthermore, we show that the algorithm guarantees optimality in a special subclass of the coverage problem. Through the experiments, we illustrate that the proposed algorithm outperforms state-of-the-art benchmark algorithms. Additionally, runtime results confirm the scalability of the algorithm, consistent with theory.
@@ Appendix A Proof of Theorem 5 (convergence)
For each iteration, Algorithm 2 will produce a new state $(x^{\prime},\mathcal{P}^{\prime})$ from an old state $(x^{\prime},\mathcal{P}^{\prime})$. Note that after step a,
$\begin{array}[]{l}M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})=\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime}).\end{array}$ (25)
Also, after step b,
$\begin{array}[]{l}M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})=\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}}).\end{array}$ (28)
@@
In case of step a, if $i_{\min}(x,\mathcal{P})\in\mathcal{N}_{i}$,
$\begin{array}[]{l}\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})-u_{\min}(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\setminus\{i_{\min}(x,\mathcal{P})\}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\setminus\{i_{\min}(x,\mathcal{P}),i_{\max}^{+}(x,\mathcal{P})\}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}}).\end{array}$ (44)
@@
$\begin{array}[]{l}\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x^{\prime},\mathcal{P}^{\prime})\\ =M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}y_{1},\ldots,y_{k}\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y_{1},\ldots,y_{k};\mathcal{P}_{n_{i}})\\ \geq\underset{\begin{subarray}{c}y_{1},\ldots,y_{k-2}\in\mathcal{P}_{n_{i}}\setminus\{i_{\min}(x,\mathcal{P}),i_{\max}^{+}(x,\mathcal{P})\}\end{subarray}}{\max}G(y_{1},\ldots,y_{k-2};\mathcal{P}_{n_{i}})+\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ \geq\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\setminus\{i_{\min}(x,\mathcal{P}),i_{\max}^{+}(x,\mathcal{P})\}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ =\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})-u_{\min}(x,\mathcal{P}).\end{array}$ (58)
If $i_{\min}(x,\mathcal{P})\not\in\mathcal{N}_{i}$, since $M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\leq u_{\min}(x,\mathcal{P})$,
$\begin{array}[]{ll}\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x^{\prime},\mathcal{P}^{\prime})&=M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\\ &\geq M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-u_{\min}(x,\mathcal{P}).\end{array}$ (62)
Since $i_{\max}^{+}\in\mathcal{N}_{i}$, $\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})=\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+M_{1}(x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})$ means a sum of utilities when $n_{i}+1$ agents share $\mathcal{N}_{i}$. Then,
$\begin{array}[]{l}M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})\geq\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P}).\end{array}$ (65)
@@
$\begin{array}[]{l}\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x^{\prime},\mathcal{P}^{\prime})\geq\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})-u_{\min}(x,\mathcal{P}).\end{array}$ (69)
In both cases, we have
@@
In case of step b, $i_{\min}(x,\mathcal{P})\not\in\mathcal{N}_{i}$ and $M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})>u_{\min}(x,\mathcal{P})$.
First, step b computes $(x^{\prime},\mathcal{P}^{\prime})$ as if an imaginary agent $l$ is added into $\mathcal{P}_{n_{i}}$, especially at $\mathcal{P}^{\prime}_{i_{+}}$. Then,
$\begin{array}[]{l}\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})=M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}})\end{array}$ (79)
@@
$\begin{array}[]{ll}M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}})\\ =M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})\\ \geq M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\\ >u_{\min}(x,\mathcal{P}).\end{array}$ (85)
Meanwhile, since $i_{\max}^{+}\in\mathcal{N}_{i}$,
$\begin{array}[]{l}\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+V(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+M_{1}(x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ \leq M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}}).\end{array}$ (93)
Then, since $i_{\min}(x,\mathcal{P})\not\in\mathcal{N}_{i}$,

### 08 Apr 2024 v3

@@ 3.3. Special Case
Even though the problem itself is no longer difficult in that case, it is noteworthy that the strategy employed in our algorithm guarantees optimality.
3.4. Communication Requirements
As shown in Theorem 4, the algorithm achieves the approximation ratio of 2, by communicating the minimum degree of information, $u_{\min}$, $i_{\max}^{+}$ and $x_{i_{\min}}$. The agent with the maximal gain from an additional agent in her region $i_{\max}^{+}$ is used to focus on the area to be reallocated in Algorithm 1. It is convenient to use it, but we note that it is not strictly required to make the algorithm work (see for instance Marden (2016)). The minimum utility, on the other hand, $u_{\min}$ is the key to proving the approximation ratio based on Equation 16. The location $x_{i_{\min}}$ is required to extend the algorithm for 1-dimensional setting in Marden (2016) to more general settings.
@@ 3.4. Communication Requirements
As for the communication complexity of the algorithm, we start by bounding the convergence rate, which is the number of iterations before convergence. Let $d_{\max}$ be the upper bound of the distance between any two points in the environment $C$, and $\Delta v$ be the resolution limit of the weight $v_{c}$. In the case of discrete settings, the potential function $\phi(x,\mathcal{P})$ consists of the elemental term $v_{c}\leavevmode\nobreak\ g(|x_{i}-c|)$ and then the improvement in the potential function for each iteration is lower bounded by $\epsilon=\Delta v\cdot g(d_{\max})$. (In the case of continuous settings, we can regard $\epsilon$ as the agents’ resolution limit of utility). Note that the convergence of the algorithm is guaranteed by Theorem 4. For each iteration, Algorithm 1 checks if the potential function can be improved or not. Before convergence, at least one out of $n$ agents improves the potential function. Then the convergence rate $\alpha$ is bounded as $\alpha\leq n[\phi(x^{*},\mathcal{P}^{*})-\phi(x,\mathcal{P})]/\epsilon$, where $(x,\mathcal{P})$ and $(x^{*},\mathcal{P}^{*})$ are the initial allocation and the allocation after convergence, respectively.
Furthermore, for each iteration, the agents build the communication tree, share the global information, share their private information to compute a neighborhood optimum, and finally share the neighborhood optimum solution. Depending on the exact algorithm used to build the communication tree, the communication burden may vary, but it can be done in $\mathcal{O}(n^{2})$. The communications to share the global information requires at most $\mathcal{O}(n^{2})$ messages.
@@ 4. Experiments
In order to validate the practical efficiency and scalability of our approach, we ran several simulations. First, we evaluate the efficiency with small graphs, then we evaluate the scalability with larger graphs.
@@
The neighborhood optimal approach proposed in Section 3.1, coined in the following as NBO, is compared to the following algorithms:
@@
  • (SOTA) the algorithm of Sadeghi et al. Sadeghi et al. (2020). In a nutshell, for a given agent $i$, the algorithm first tries to perform individual moves to maximize the (local) social welfare of her neighborhood, as defined by the Voronoi partitioning. If no such move is improving, it considers coordinated moves with a single other agent $j$, in the sense that $i$ would move and $j$ would take the place of agent $i$. The algorithm first considers neighbors, and then (via neighborhood communication), may consider agents further away. However, these coordinated moves only involve two agents at most.
  • (CGR) the centralized greedy algorithm that allocates agents one by one starting from the empty environment. Recall that this algorithm guarantees a performance of $1-1/e\approx$ 63% of the optimal solution.
We evaluate the performance of the algorithms above with different shapes of the environments (Figure 7). In addition to these shapes, we also use the small bridge setting and OR library dataset shown in Yun and Rus (2014). More details about the shapes and the way instances are generated are available in the supplementary material.
@@ 4.1. Efficiency
First, we evaluate the efficiency of the proposed neighborhood optimum algorithm, by comparing its solution with the optimal one and with solutions returned by VVP, SOTA and CGR. The efficiency is measured as the ratio $G(x)/G(x^{*})$ where $x$ is a solution of the algorithm and $x^{*}$ is an optimal solution. Since finding an optimal solution is NP-hard, the simulation is conducted on a 1D chain with $|C|=20,|\mathcal{C}_{+}|=10$, and $N=5$.
In that case, the proposed algorithm outperforms both distributed algorithms (VVP and SOTA) and improves the efficiency by about 21.7% compared to SOTA (Figure 8). The efficiency ratio of NBO is 90% which is much better than the theoretical approximation ratio of 2. We examine the detail with an illustrative example in the supplementary material.
Due to the NP-hardness of the problem, it is difficult to compare with optimal solutions in larger settings. For that reason, we instead compute the efficiency ratio with the centralized greedy (CGR) as a practical baseline. Table 1 summarizes all the results obtained. The quantitative comparison of the efficiency shows similar results as in Figure 8 that the proposed method outperforms the benchmarks in all cases. Note that NBO even sometimes outperforms CGR in some instances, despite the following two disadvantages of NBO compared to CGR: (i) NBO uses only limited information while CGR uses full information, (ii) NBO starts from a random initial state, which is less favorable compared to CGR’s iterative approach from the empty environment.
@@
Table 1. Comparison on different shapes: Efficiency ratio $G(x)/G(x^{CGR})$ where $x$ is a solution of each algorithm and $x^{CGR}$ is a solution of CGR. Each cell shows mean ± standard deviation. Larger values mean better results.
@@ 4.2. Scalability
We finally evaluate the scalability of the proposed algorithm by varying the size of the space $|C|$ and the number of agents $n$ (we could run experiments with more than 150 agents, the details can be found in the supplementary material). The results show that:
  • The runtime grows with the size $|C|$, as expected. On the largest instances of size $|C|=192$, with 20 agents, the runtime is above 3 hours.
  • The runtime decreases when $n$ increases because the size of partitions $|\mathcal{P}_{ij}|$ also decreases, and thus the opportunities of improvement are more severely constrained.
These results demonstrate the applicability of the proposed algorithm to real-world medium-scale problems. For larger-scale settings, further improvements will be required to make the approach viable. It should be emphasized though that the anytime nature of the algorithm makes it relevant even with a limited time budget. Further details about the experiments are deferred to the supplementary material.
@@ 5. Conclusion
This paper extends the approach of Marden (2016) to general dimensional settings for discrete environments, ensuring convergence to a neighborhood optimum solution, even in challenging non-convex scenarios, with an approximation ratio of 2. The communication requirements involve disseminating the value and position of the minimum utility agent, capturing an approximation ratio that surpasses other methods presented in Sadeghi et al. (2020); Durham et al. (2011). Interestingly, this minimal level of informational dissemination recaptures the best achievable approximation ratio of 2 for Nash equilibria, which requires all agents to have full information about the environment (e.g., choices of all agents, utility associated with all feasible choices, etc.). While more communication demanding than simple best responses based on Voronoi partitioning, the informational dissemination requirements are manageable, enabling local decision-making rules. Furthermore, our algorithm guarantees optimality in a special subclass of the coverage problem and outperforms state-of-the-art benchmark algorithms in experiments. Future research will focus on improving the algorithm’s efficiency by minimizing information transmission or enhancing experimental results. Implementing this algorithm on real robotic systems is another important avenue for further exploration.
{acks}
This work is partially supported by ONR grant #N00014-20-1-2359 and AFOSR grants #FA9550-20-1-0054 and #FA9550-21-1-0203.
@@ Appendix A Proof of Theorem 4 (convergence)
Here, a pair of neighboring agents $S=\{i,j\}$ is randomly chosen to update their joint actions. For simplicity, we assume $|S|=2$ to explain, but the following proofs hold even for the case of $|S|>2$. Without loss of generality, we assume that $i$ is either the root of the minimum utility tree or closer than $j$ according to the communication tree. For each iteration, Algorithm 1 will produce a new state $(x^{\prime},\mathcal{P}^{\prime})$ from an old state $(x,\mathcal{P})$. Note that after step a, where agents in $S$ reallocate $\mathcal{P}_{S}$ by themselves,
$\begin{array}[]{l}M_{2}(\emptyset,\mathcal{P}_{S})=\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime}).\end{array}$ (21)
Also, after step b, where agents in $S$ try to make a space in $\mathcal{P}_{S}$ to accommodate another agent,
$\begin{array}[]{l}M_{3}(\emptyset,\mathcal{P}_{S})=\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}}).\end{array}$ (24)
@@
In case of step a, if $i_{\min}(x,\mathcal{P})\in S$,
$\begin{array}[]{lcl}\underset{\begin{subarray}{c}j\in S\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})-u_{\min}(x,\mathcal{P})&=&\underset{\begin{subarray}{c}j\in S\setminus\{i_{\min}(x,\mathcal{P})\}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})\\ &=&\underset{\begin{subarray}{c}j\in S\setminus\{i_{\min}(x,\mathcal{P}),i_{\max}^{+}(x,\mathcal{P})\}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+u_{i_{\max}^{+}}(x,\mathcal{P})\\ &&+\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})-u_{i_{\max}^{+}}(x,\mathcal{P})\\ &=&\underset{\begin{subarray}{c}j\in S\setminus\{i_{\min}(x,\mathcal{P}),i_{\max}^{+}(x,\mathcal{P})\}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}}).\end{array}$ (43)
@@
$\begin{array}[]{l}\underset{\begin{subarray}{c}j\in S\end{subarray}}{\sum}u_{j}(x^{\prime},\mathcal{P}^{\prime})\\ =M_{2}(\emptyset,\mathcal{P}_{S})\\ =\underset{\begin{subarray}{c}y_{1},\ldots,y_{k}\in\mathcal{P}_{S}\end{subarray}}{\max}G(y_{1},\ldots,y_{k};\mathcal{P}_{S})\\ \geq\underset{\begin{subarray}{c}y_{1},\ldots,y_{k-2}\in\mathcal{P}_{S}\setminus\{i_{\min}(x,\mathcal{P}),i_{\max}^{+}(x,\mathcal{P})\}\end{subarray}}{\max}G(y_{1},\ldots,y_{k-2};\mathcal{P}_{S})+\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ \geq\underset{\begin{subarray}{c}j\in S\setminus\{i_{\min}(x,\mathcal{P}),i_{\max}^{+}(x,\mathcal{P})\}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ =\underset{\begin{subarray}{c}j\in S\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})-u_{\min}(x,\mathcal{P}).\end{array}$ (57)
If $i_{\min}(x,\mathcal{P})\not\in S$, since $M_{3}(\emptyset,\mathcal{P}_{S})-M_{2}(\emptyset,\mathcal{P}_{S})\leq u_{\min}(x,\mathcal{P})$,
$\begin{array}[]{ll}\underset{\begin{subarray}{c}j\in S\end{subarray}}{\sum}u_{j}(x^{\prime},\mathcal{P}^{\prime})&=M_{2}(\emptyset,\mathcal{P}_{S})\\ &\geq M_{3}(\emptyset,\mathcal{P}_{S})-u_{\min}(x,\mathcal{P}).\end{array}$ (61)
Since $i_{\max}^{+}\in S$, $\underset{\begin{subarray}{c}j\in S\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})=\underset{\begin{subarray}{c}j\in S\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+M_{1}(x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})$ means a sum of utilities when $3$ agents share $S$. Then,
$\begin{array}[]{l}M_{3}(\emptyset,\mathcal{P}_{S})\geq\underset{\begin{subarray}{c}j\in S\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P}).\end{array}$ (64)
@@
$\begin{array}[]{l}\underset{\begin{subarray}{c}j\in S\end{subarray}}{\sum}u_{j}(x^{\prime},\mathcal{P}^{\prime})\geq\underset{\begin{subarray}{c}j\in S\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})-u_{\min}(x,\mathcal{P}).\end{array}$ (68)
Note that $V(x,\mathcal{P})>u_{\min}(x,\mathcal{P})$ in $Z_{1}$. Therefore, in both cases ($i_{\min}(x,\mathcal{P})\in S$ and $i_{\min}(x,\mathcal{P})\not\in S$), we have
@@
In case of step b, $i_{\min}(x,\mathcal{P})\not\in S$ and $M_{3}(\emptyset,\mathcal{P}_{S})-M_{2}(\emptyset,\mathcal{P}_{S})>u_{\min}(x,\mathcal{P})$.
First, step b computes $(x^{\prime},\mathcal{P}^{\prime})$ as if an imaginary agent $l$ is added into $\mathcal{P}_{S}$, especially at $\mathcal{P}^{\prime}_{i_{+}}$. Then,
$\begin{array}[]{l}\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})=M_{3}(\emptyset,\mathcal{P}_{S})-M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}})\end{array}$ (78)
@@
$\begin{array}[]{ll}M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}})\\ =M_{3}(\emptyset,\mathcal{P}_{S})-\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})\\ \geq M_{3}(\emptyset,\mathcal{P}_{S})-M_{2}(\emptyset,\mathcal{P}_{S})\\ >u_{\min}(x,\mathcal{P}).\end{array}$ (84)
Meanwhile, since $i_{\max}^{+}\in S$,
$\begin{array}[]{l}\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+V(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+M_{1}(x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ \leq M_{3}(\emptyset,\mathcal{P}_{S})\\ =\underset{\begin{subarray}{c}i\in S\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}}).\end{array}$ (92)
Then, since $i_{\min}(x,\mathcal{P})\not\in S$,